The combustion of liquid methanol has a delta H of -1278 kJ how much energy will you get from burning 245.0 g of methanol in excess oxygen. 2CH3OH(l) + 3O2(g) - 2CO2(g) + 4H2O(g)
Answer #1

Molar mass of CH3OH=12+3+16+1=32

Number of moles in 245 gram of methanol=245/32=7.65625

Since ,in a combustion reaction energy is of one mole only.

So energy got from 245g methanol

=7.65625×-1278= -9784.6875 kJ